# Anonymization by Design of Language Modeling

### v1

@@ Abstract.
Rapid advances in Natural Language Processing (NLP) have revolutionized many fields, including healthcare. However, these advances raise significant privacy concerns, especially when models specialized on sensitive data can memorize and then expose and regurgitate confidential information. This paper presents a privacy-by-design language modeling approach to address the problem of language models anonymization, and thus promote their sharing. Specifically, we propose both a Masking Language Modeling (MLM) methodology to specialize a BERT-like language model, and a Causal Language Modeling (CLM) methodology to specialize a GPT-like model that avoids the model from memorizing direct and indirect identifying information present in the training data. We have comprehensively evaluated our approaches using medical datasets and compared them against different baselines. Our results indicate that by avoiding memorizing both direct and indirect identifiers during model specialization, our masking and causal language modeling schemes offer the best tradeoff for maintaining high privacy while retaining high utility.
@@ 1. Introduction
The healthcare sector is a major generator of sensitive data collected from different sources. The exploitation of this valuable data presents many promises such as improving the quality of care, acquiring a better knowledge of the healthcare system, identifying disease risk factors, aiding in diagnosis, personalized care to name a few. A large amount of this data is unstructured text documents in the form of medical reports. With the rise of Machine Learning (ML) and the advent of Natural Language Processing (NLP), language models are increasingly used to automate the processing of these medical reports (Chen et al., 2022; Yeo et al., 2022; Wei et al., 2022; Goyal et al., 2024).
Patient medical records are extremely sensitive and private data. Their use and dissemination are therefore subject to numerous regulations such as HIPAA, for the USA, or GDPR for Europe. In these regulations, one of the main prerequisites for the dissemination of medical data is to remove any element that allows a patient to be directly identified (i.e., de-identification or pseudonymization) or indirectly (i.e., anonymization).
@@
Recent advances in NLP (Singh and Mahmood, 2021) based on neural networks have democratized their use. Since the advent of ChatGPT, NLP models are not limited to text generation and can include several tasks, including classification and Named Entity Recognition (NER), which therefore allows for de-identification of free texts. For example, Johnson et al. proposed to use a neural network based on a BERT architecture (et al., 2020) to detect a number of identifying elements in medical documents. More recently, different hospitals have also explored the feasibility of using NLP models to automatically pseudonymize text documents (i.e., hide specific direct identifiers named Personally Identifiable Information (PII)) from their clinical data warehouse (Tannier et al., 2023; Richard et al., 2023). In these approaches, the BERT model is fine-tuned with the medical reports from the hospital (in order to specialize and well understand the reports generated by the local practitioners) before training a Named Entity Recognition on a set of Personally Identifiable Information that directly identify patients.
@@
However, pseudonymization is not sufficient for text anonymization (Pilán et al., 2022). Indeed, the identity of an individual can be found from indirect identifying information. Therefore, these indirect identifiers must also be protected (e.g., removed) in order to properly anonymize texts. This operation cannot rely on a classification or NER task, because there is no a priori information to identify these indirect identifiers, it depends on the context and the dataset considered. To our knowledge, no NLP model training methodology has been proposed for taking into account indirectly re-identifying information in the protection of medical reports or documents from other domains.
@@
The need for information sharing in the medical field goes beyond data sharing and now concerns learning models. As language models specialized on clinical reports improves their performance for downstream clinical NLP tasks (Huang et al., 2020; Labrak et al., 2023; Dura et al., 2022), some hospitals or healthcare centers also want to be able to share models specialized and fine-tuned on their own data with other centers. For example, a hospital specialized in oncology may train a classification model with its own data and would like to share this model with other hospitals. In this context, the model specialized with the medical reports of the first hospital must not be able to regurgitate sensitive data used during its specialization (or return any other information allowing to deduce them) when it will be used by the second hospital. The attack surface on models trained on personal and highly sensitive data is still poorly understood (Weidinger et al., 2022; Lehman et al., 2021; Duprieu and Berkouk, 2024). A number of threats are related to the memorization and possible leakage of sensitive information used during model training. There are some known privacy vulnerabilities involving training data associated with learning models, such as memorization, data reconstruction, and membership inference (i.e., identifying elements used during the training).
@@
Memorization of information by a model is not a problem in itself, it is even necessary for learning. However, this memorization becomes a problem when the training information is not generalized enough by the model which reproduces large portions of training data verbatim or discloses some sensitive information. And more specifically, the risk to privacy is increased when this problematic memorization concerns identifying information. Indeed, memorizing identifiers that leads to the regurgitation of these identifiers can clearly be used to infer that the information of a specific individual was used during the training of the model (i.e., membership inference). Thus, in order to share a model without risk of regurgitation of personal information, it is necessary to ensure that a model has not memorized both direct and indirect identifiers. However, as far as we know, no methodology has been proposed to help the different actors, especially hospitals, to ”anonymize” their learning models before sharing them.
To reduce privacy risks, mitigation techniques have been proposed such as learning with differential privacy (Abadi et al., 2016) (DP) guarantees or pruning strategy (Wang et al., 2022). However, these mitigation techniques only increase the probability of the adversary to make a wrong decision or inference. Moreover, these countermeasures significantly degrade the accuracy of the model, making them unusable in practice. Finally, these techniques do not have a specific treatment for directly or indirectly identifying data.
@@
In this paper, we propose privacy-by-design language modeling approaches to address the problem of anonymizing language models. Specifically, we propose both a masking language modeling (MLM) methodology to specialize a BERT-like language model, and a causal language modeling (CLM) methodology to specialize a GPT-like model that avoids the model from memorizing direct and indirect identifying information. To achieve this goal, our fine-tuning approaches named PPmlm-bert and PPclm-gpt (for a BERT and GPT model, respectively), first identify directly and indirectly identifying information in the data corpus (e.g., medical reports). While direct identification words are identified by exploiting name entity recognition, indirect identification words are identified as words used only by a single patient. This ensures that the memorized words have been used in the documents of at least two individuals (i.e., thus providing k-anonymity, note that the value of k can be parameterized in our solution). Then, the specialization operation (i.e., the random masking or the next-word prediction for BERT and GPT models, respectively) avoids using directly and indirectly identifying words. Therefore, the model is specialized on the dataset without ever learning identifying words, which then drastically reduces the risk of regurgitation or inference of personal information during its use.
We have comprehensively evaluated our approaches using datasets of english and french medical reports and considering a large number of baselines for comparison. We first illustrate the risk of identifying patients whose reports were used for model specialization, thus motivating our approaches. We then show that ignoring identifier words in training (i.e., direct and indirect identifiers) improves privacy by reducing the ability of models to predict these identifier words while maintaining good ability to predict other non-identifier words (aka the utility), or the ability of the model to train a downstream classification task.
Our contributions are as follows:
  1. (1)
    we illustrate the risk of indirectly identifying data regurgitation (i.e., privacy leakage of training data) of NLP models following its specialization, motivating our anonymization solution for model specialization;
  2. (2)
    we provide a privacy-by-design approach for masked and causal language modeling that ensures that direct and indirect identifiers are not memorized during specialization;
  3. (3)
    we comprehensively evaluate our approach and compare it to classical fine-tuning of BERT and GPT models on medical corpora, as well as to specialization with de-identified datasets.
@@
The paper is organized as follows. Section 2 reviews background and related work while Section 3 defines the problem statement. Section 4 describes our privacy-preserving language modeling for both descriptive and generative models. Section 5 then presents the experimental setup while the evaluation results are reported in Section 6. Finally, we discuss limitations and future works Section 7 before concluding Section 8.
@@ 2.3. Privacy Leakages
A central question in this context concerns the extent to which language models memorize their training data (Carlini et al., 2019, 2023; Nasr et al., 2023; Zhang et al., 2021; Schwarzschild et al., 2024; Speicher et al., 2024). However, defining memorization for language models is challenging, and many existing definitions and notions have been proposed depending on whether the memorization concerns copyrighted content or personal and sensitive content. In relation to privacy, we can notably cite extractable memorization (Section 2.3.2) and counterfactual memorization (Section 2.3.1). This memorization has a direct impact on privacy risks. For example, we illustrate Section 6.1 the exploitation of this memorization to perform a membership inference (Section 2.3.3).
2.3.1. Counterfactual memorization
ML models are supposed to learn general information. Rare data, on the other hand, is not supposed to be memorized. A model that memorizes rare data (sometimes called outliers) not only has a negative impact on utility but also on privacy. Indeed, the more we learn about a small subset of individuals, the greater the information leakage because we can more easily trace that information back to its source. For example, we expect chatGPT to know Harry Potter’s address (which can be found on many pages online) but not the reader’s address (which should be nonexistent or at least difficult to find online).
It turns out that it is possible to measure this unwanted memorization, called counterfactual memorization in (Zhang et al., 2021). To do this on any data, you need to compare the performance of a model trained on a dataset with that data, with a second model trained without it. This is computationally expensive for each data set, so counterfactual memorization is actually calculated with an empirical expectation: we create multiple copies of the dataset and train many models on different subsets. Each data set will have models it was trained on and models it was not trained on. We can then calculate the expected memorization:
$\textrm{mem}(x)=E_{x\in D}(\textrm{score}(M_{D},x))-E_{x\notin D^{\prime}}(\textrm{score}(M_{D^{\prime}},x)),$
where $\textrm{score}(M_{D},x)$ is the score for $x$ of the model trained with the dataset $D$. The two terms will cancel for common data (removing them has no impact) but can give a high difference for rare data. Data points with memorability above a certain threshold will be considered at risk.
Counterfactual memorization aims to separate memorization from generalization and requires training the model multiple times, which is a limitation in practice given the cost of training a model.
@@ 2.3.3. Membership inference
Membership inference attack (Carlini et al., 2022) (MIA) is a more common inference attack in machine learning, which aims to infer whether a specific data was used in the training data of a target model. This can be a problem, for example, if a hospital has trained a model to detect cancer and you learn that your colleague’s data was used in the training. You will have indirectly learned that she probably has cancer.
There are different techniques that can be used to perform a MIA attack. One of them is to use shadow models (Shokri et al., 2017). Shadow models are trained to mimic the behavior of the target model on an auxiliary dataset with a similar distribution to the original. An adversarial model (i.e., a classifier) is then trained to infer membership from these shadow models.
Although membership inference attacks have been used to quantify memorization risks (Jagannatha et al., 2021; Mireshghallah et al., 2022; Wang et al., 2022), it may not be applicable in a practical scenario due to the high dimensional input space. By ignoring and not memorizing the direct and indirect identifiers of the training data during model specialization, our privacy-by-design scheme avoids membership inferences targeted on identifiers.
@@ 2.4. Mitigation strategies
where $x,y$ represent data and its label, $M_{D}$ a model trained on dataset D and $\epsilon$ the  privacy budget. The lower $\epsilon$ is, the more private the model is.
@@
The most popular method to apply DP in machine learning is DP-SGD: Differentially-Private Stochastic Gradient Descent (Abadi et al., 2016). The idea is to apply DP during the training phase by clipping gradient updates and adding centered noise at each step. DP is known to significantly decrease the accuracy of the model (Jagannatha et al., 2021) and privacy budget management is difficult.
3. Problem statement
@@
As the training data contained personal information, our goal is to avoid that the model memorizes any information that can be used to identify individuals which information was used during model training. This information concerns both directly and indirectly identifying data. Directly identifiable data includes information associated with a natural person such as first and last name, a telephone or license plate number, an identifier such as a social security number, a postal or email address. Indirectly identifiers, in turn, is information that is uniquely used in an individual’s data compared to the rest of the population. For example, if the term vertebra is used only in a patient’s data, this term is indirectly identifying compared to the rest of the corpus. In other words, the individual is distinguishable from others with the use of this term. Figure 3 depicts a bipartite graph which relates the people represented in the data corpus to the words used across the entire dataset. More formally, it is a Bipartite Graph $G=(V,E)$ where the vertices of this graph are the users and words of the dataset $V=(U\bigcup I)$. The use of a word in a user’s data is represented in $G$ by an edge $(u,i)\in U\texttimes I$. In this example, we can see that the words $w1$ and $w4$ are only used in the data of individual $u1$. Note that a directly identifying word can also be a word used only in an individual’s data, this is the typical case of a last name or a particular identifier.
@@
Figure 3. By constructing a bipartite graph between individuals and the words used in their documents, we can easily identify indirectly identifying words (i.e., words pointed to only by one individual).
@@
These directly and indirectly identifying data can pose a problem if they are memorized by the model following training. Indeed, an identifying word memorized by the model which would be then proposed in the test phase would betray its use during training. As depicted Figure 4, the model can be made to memorize the context information and the information that was used as a mask in the supervised training phase [^1] We discuss the memorization related to the context in Section 7 and show that its privacy impact is very limited compared to memorization of words used in masks.. In this example, the fact that the model proposes the word $w4$ in the testing phase, mean s that an adversary could deduce that the data of user $u1$ was used during training because this word only appears in the data of this individual.

### v2

# Towards the Anonymization of the Language Modeling

@@ Abstract.
Rapid advances in Natural Language Processing (NLP) have revolutionized many fields, including healthcare. However, these advances raise significant privacy concerns, especially when pre-trained models fine-tuned and specialized on sensitive data can memorize and then expose and regurgitate personal information. This paper presents a privacy-preserving language modeling approach to address the problem of language models anonymization, and thus promote their sharing. Specifically, we propose both a Masking Language Modeling (MLM) methodology to specialize a BERT-like language model, and a Causal Language Modeling (CLM) methodology to specialize a GPT-like model that avoids the model from memorizing direct and indirect identifying information present in the training data. We have comprehensively evaluated our approaches using a medical dataset and compared them against different baselines. Our results indicate that by avoiding memorizing both direct and indirect identifiers during model specialization, our masking and causal language modeling schemes offer a good tradeoff for maintaining high privacy while retaining high utility.
@@ 1. Introduction
The healthcare sector is a major generator of sensitive data collected from different sources. The exploitation of this valuable data presents many promises such as improving the quality of care, acquiring a better knowledge of the healthcare system, identifying disease risk factors, aiding in diagnosis, personalized care to name a few. A large amount of this data is unstructured text documents in the form of medical reports. With the rise of Machine Learning (ML) and the advent of Natural Language Processing (NLP), language models are increasingly used to automate the processing of these medical reports (Chen et al., 2022a; Yeo et al., 2022; Wei et al., 2022; Goyal et al., 2024).
@@
Patient medical records are extremely sensitive and private data. Their use and dissemination are therefore subject to numerous regulations such as HIPAA (Health Insurance Portability and Accountability Act), for the USA, or GDPR (General Data Protection Regulation) for Europe. In these regulations, one of the main prerequisites for the dissemination of medical data is to remove any element that allows a patient to be directly identified (i.e., de-identification or pseudonymization) or indirectly (i.e., anonymization). Recent advances in NLP (Singh and Mahmood, 2021) based on neural networks have democratized their use. Since the advent of ChatGPT, NLP models are not limited to text generation and can include several tasks, such as classification tasks or Named Entity Recognition (NER) which therefore allows for de-identification of free texts if it is trained to recognize identifiers. For example, Johnson et al. proposed to use a neural network based on a BERT architecture (et al., 2020) to detect a number of identifying elements in medical documents. More recently, different hospitals have also explored the feasibility of using NLP models to automatically pseudonymize text documents, means hiding specific direct identifiers named Personally Identifiable Information (PII) from their clinical data warehouse (Tannier et al., 2023; Richard et al., 2023). The set of PII to consider is recommended by Data Protection Authorities (DPA)  [^1] For instance, in Section 5.2.1.1 of the reference document relating to the processing of personal data implemented for the purpose of creating data warehouses in the health sector from the French DPA (i.e., CNIL): https://www.cnil.fr/sites/cnil/files/atoms/files/referentiel_entrepot.pdf and includes names, hospitals, addresses, birth dates, patient numbers for instance.
@@
Although a large number of works only exploit NER to pseudonymize documents (Speicher et al., 2024), pseudonymization is not sufficient for text anonymization (Pilán et al., 2022). Indeed, a link between an individual and their data can be found from indirect identifying information. Therefore, these indirect identifiers must also be protected (e.g., removed) in order to better anonymize texts. This operation cannot rely on a classification or NER task as there is no a priori information to identify these indirect identifiers, it depends on the context and the dataset considered. To our knowledge, no NLP model training methodology has been proposed for taking into account indirectly re-identifying information in the protection of medical reports or documents from other domains.
@@
The need for information sharing in the medical field goes beyond data sharing and now concerns learning models. As language models specialized on clinical reports improves their performance for downstream clinical NLP tasks (Huang et al., 2020; Labrak et al., 2023; Dura et al., 2022), some hospitals or healthcare centers also want to be able to share pre-trained models specialized and fine-tuned on their own data with other medical centers. For example, a hospital specialized in oncology may train a classification model with its own data and would like to share this model with other hospitals. In this context, the model specialized with the medical reports of the first hospital must not be able to regurgitate sensitive data used during its specialization (or return any other information allowing to deduce them) when it will be used by the second hospital. The attack surface on models trained on personal and highly sensitive data is still poorly understood (Weidinger et al., 2022; Lehman et al., 2021; Duprieu and Berkouk, 2024). A number of threats are related to the memorization and possible leakage of sensitive information used during model training. There are some known privacy vulnerabilities involving training data associated with learning models, such as memorization, data reconstruction, and membership inference (i.e., identifying elements used during the training or the fine-tuning).
@@
Memorization of information by a model is not a problem in itself, it is even necessary for learning. However, this memorization becomes a problem when the training information is not generalized enough by the model which reproduces large portions of training data verbatim or discloses some sensitive information. And more specifically, the risk to privacy is increased when this problematic memorization concerns identifying information. Indeed, memorizing identifiers that leads to the regurgitation of these identifiers can clearly be used to infer that the information of a specific individual was used during the training of the model (i.e., membership inference). Thus, in order to share a model without risk of regurgitation of personal information, it is necessary to ensure that a model has not memorized both direct and indirect identifiers. However, as far as we know, no methodology has been proposed to help model designers, especially hospitals, to ”anonymize” their learning models before sharing them.
As part of investigations into the training of AI models, the Irish DPA has asked the European Data Protection Board (EDPB) to issue an opinion on the development and deployment of AI models  [^2] https://www.edpb.europa.eu/news/news/2024/edpb-opinion-ai-models-gdpr-principles-support-responsible-ai_en in light of the GDPR. More precisely, the opinion addresses the issues of AI models trained with personal data, regardless if the model is designed to provide information about individuals whose data is included in their training sets or not. The opinion therefore considers that an AI model trained on personal data cannot be considered anonymous in all cases, and that in most cases, an analysis will have to be conducted by the data controller to demonstrate that such a model is anonymous. To be considered anonymous, the likelihood of regurgitation of personal information must be insignificant taking into account all the means reasonably likely to be used.
To reduce privacy risks, mitigation techniques have been proposed such as learning with Differential Privacy (DP) guarantees (Abadi et al., 2016) or pruning strategy (Wang et al., 2022). While these mitigation techniques increase the probability of the adversary to make a wrong decision or inference, these countermeasures also significantly degrade the accuracy of the model, making them difficult to use in practice. Finally, these techniques are data agnostic and do not treat identifying data differently from other data. DP was originally designed as a countermeasure to membership inference attacks (MIAs). MIAs aim to predict whether a particular record belongs to a given model’s training dataset. This analysis addresses not only model privacy auditing but also copyright infringement detection. The underlying concept of this attack relies on the fact that a model behaves differently depending on whether it is processing data that it has already seen during training or not. Most of the state-of-the-art attacks attempt to capture this change in behavior and require training multiple models, making the attack very expensive. Furthermore, the performance of existing MIAs on LLMs is more and less conclusive (Duan et al., 2024; Eichler et al., 2025; Mireshghallah et al., 2022).
@@
In this paper, we propose privacy-by-design language modeling approaches to address the problem of anonymizing language models. Specifically, we propose both a masking language modeling (MLM) methodology to specialize a pre-trained BERT-like language model, and a causal language modeling (CLM) methodology to specialize a pre-trained GPT-like model that avoids the model from memorizing direct and indirect identifying information. To achieve this goal, our fine-tuning approaches named PPmlm-bert and PPclm-gpt (for a BERT and GPT model, respectively), first identify directly and indirectly identifying information in the data corpus used for the specialization. While direct identifiers are pinpointed by exploiting a off-the-shelf NER model, indirect identification words are identified as words used only by a single patient (which goes beyond the directly identifying words). This ensures that the memorized words were used in the data of at least two individuals (i.e., thus making the k=2 individuals indistinguishable, note that the value of k can be parameterized in our solution). Then, the language modeling to specialize and fine-tune the model on the training data (i.e., the random masking or the next-word prediction for BERT and GPT models, respectively) avoids using directly and indirectly identifying words. Therefore, the risk of regurgitation or inference of personal information during the use of the model is drastically reduced.
We comprehensively evaluated our approaches using medical reports and considering a wide range of comparison baselines (e.g., the use of pseudonymized data for training, DP). We quantified the impact of our solution on the resulting model’s performance using multiple utility metrics (e.g., the ability to predict words and tokens, training a downstream classification task), as well as on privacy metrics based on the model’s ability to predict identifiers or through MIAs. We show that ignoring direct and indirect identifiers in the language modeling improves privacy while maintaining good utility. To enable reproducibility of results and appropriation of the solution, code and public data as well as other associated research artifacts are available [^3] Code for reproducibility: anonymized for submission.
To summarize our contribution, this work addresses the too-often-seen limitation that relying on NER models to detect identifying information and simply removing it from the training data is sufficient to adequately protect privacy. We believe that recognizing this limitation and proposing a solution could be useful in practice to prevent models supposedly without personal information (i.e., who have only exploited a NER) from being freely exchanged on Hugging Face.
@@
The paper is organized as follows. Section 2 reviews the background and related work while Section 3 defines the threat model. Section 4 describes our privacy-preserving language modeling for both descriptive and generative models. Section 5 then presents the experimental setup while the evaluation results are reported in Section 6. Finally, we discuss the limitations and future work Section 7 before concluding Section 8.
@@ 2.3. Privacy Leakages
A central question in this context concerns the extent to which language models memorize their training data (Carlini et al., 2019, 2023; Nasr et al., 2023; Zhang et al., 2021; Schwarzschild et al., 2024). However, defining memorization for language models is challenging, and many existing definitions and notions have been proposed depending on whether the memorization concerns copyrighted content or personal and sensitive content. In relation to privacy, we can notably cite extractable memorization (Section 2.3.1) and membership inference (Section 2.3.2).
@@ 2.3.2. Membership inference
Membership inference attack (Carlini et al., 2022) (MIA) is a more common inference attack in machine learning, which aims to infer whether a specific data was used in the training data of a target model. This can be a problem, for example, if a hospital has trained a model to detect risk of cancer recurrence and you learn that your colleague’s data was used in the training. You will have indirectly learned that she had cancer.
There are different techniques that can be used to perform a MIA attack depending on if the adversary has an access to the model parameters (i.e., white-box access), or access to a ground-truth subset of member and non-member samples. One technique consists to analyze the loss of member and non member samples (Yeom et al., 2018), another one is to use multiple shadow models (Shokri et al., 2017; Ye et al., 2022) trained to mimic the behavior of the target model on an auxiliary dataset. An adversarial model is then trained to infer membership from the loss or from shadow models. Another method (Zhang et al., 2021) is based on comparing the performance of the target model trained on a dataset with a specific input, with a second model trained without it. As ML models are supposed to learn general information, one piece of data (even rare, outlier or mislabeled samples) is not supposed to be memorized and significantly changed the model’s performance. By repeating this operation many times with different subset, it is possible to identify counterfactually memorized data.
Although membership inference attacks have been used to quantify memorization risks (Jagannatha et al., 2021; Mireshghallah et al., 2022; Wang et al., 2022; Duan et al., 2024; Aerni et al., 2024), it may not be applicable in a practical scenario due to the high dimensional input space (i.e., too expensive in terms of calculation). By ignoring and not memorizing the direct and indirect identifiers of the training data during language modeling, our privacy-by-design scheme avoids membership inferences targeted on identifiers.
Being able to audit privacy at low-cost is important and challenging (Zarifzadeh et al., 2024). Indeed, using MIAs that require training multiple models is intractable for LLMs. We measure the privacy level of our solution through a low-cost MIA attack that relies on the analysis of the model’s predictions. More precisely, the prediction of a word associated with an individual (i.e., a direct or indirect identifier) is considered a privacy leakage, as it betrays the fact that the model has seen this data during its training. This attack is consistent in the medical use case with off-the-shelf foundation models that have never seen the data used for language modeling.
@@ 2.4. Mitigation strategies
where $x,y$ represent data and its label, $M_{D}$ a model trained on dataset D and $\epsilon$ the  privacy budget. The lower $\epsilon$ is, the more private the model is. The most popular method to apply DP in machine learning is DP-SGD: Differentially-Private Stochastic Gradient Descent (Abadi et al., 2016). The idea is to apply DP during the training phase by clipping gradient updates and adding centered noise at each step. DP is known to significantly decrease the accuracy of the model (Jagannatha et al., 2021) and privacy budget management is difficult.
To reduce privacy risks, Chen et al. (Chen et al., 2022b) proposed to relax the loss in order to reduce the distinguishability between the training and testing loss distributions, and Ruizhe et al. (Chen et al., 2024) proposed a method to localize specific neurons responsible for memorizing PII in LLMs through adversarial training and then disable these neurons.
3. Threat Model
Since the training data contained personal information, the adversary’s goal is to identify the individuals whose information was used in training the model. To achieve this goal, the adversary takes advantage of direct and indirect identifying data from its auxiliary knowledge. Both direct and indirect identifiers are complementary. Direct identifiers include information associated with a natural person such as first and last name, a telephone or a social security number. In case of medical data, a list of direct identifiers to cosider has been recommended by the Data Protection Authority [^4] The list of direct identifiers recommended by the french DPA: https://www.cnil.fr/sites/cnil/files/atoms/files/referentiel_entrepot.pdf.. However, direct identifiers fail to detect all terms that can lead to distinguishing an individual. This distinguishing nature is captured by indirect identifiers. Indirect identifiers is information that is uniquely used in an individual’s data compared to the rest of the population. For example, if the term vertebra is used only in a patient’s data, this term is indirectly identifying compared to the rest of the corpus. In other words, the individual is distinguishable from others with the use of this term. This distinguishable nature depends on the adversary’s auxiliary knowledge. We consider an adversary which has collected information about individuals. This information can be user profiles from Web pages or mobile applications, medical reports, or any other text-type information describing an individual. These texts can be plain or pseudonymized and come from web scraping or from a legally or illegally acquired dataset. In a pseudonymized dataset, directly identifying information will be removed, but indirect identifiers will still be present. The number of direct and indirect identifiers will vary depending on the amount of information the attacker is able to retrieve on each individual, as well as the similarity of the retrieved information and the information used during training of the target model.
@@ 3. Threat Model
While the direct identifiers can be detected by a off-the-shelf NER model  [^5] EDS-Pseudo trained from Clinical Data Warehouse for instance - https://github.com/aphp/eds-pseudo, indirect identifiers have to be detected. To achieve this goal, an analysis of the data corpus can be carried out. Figure 3 depicts a bipartite graph which relates the individuals represented in the data corpus to the words used across the entire dataset. More formally, it is a Bipartite Graph $G=(V,E)$ where the vertices of this graph are the users and words of the dataset $V=(U\bigcup I)$. The use of a word in a user’s data is represented in $G$ by an edge $(u,i)\in U\texttimes I$. In this example, we can see that the word $w1$ and $w4$ are only used in the data of individual $u1$. Note that a direct identifier can also be a word used only in an individual’s data (i.e., an indirect identifier), this is the typical case of a last name.
@@
Figure 3. Detection of the identifiers: 1) a off-the-shelf NER model can detect the direct identifiers while by constructing a bipartite graph between individuals and the words used in their documents, we can easily identify indirectly identifiers (i.e., words pointed to only by one individual).
@@
These identifying data (direct and indirect) can pose a confidentiality problem if they are memorized by the model after the training phase. Indeed, an identifier memorized by the model and proposed (i.e., predicted in a masking phase) or regurgitated (i.e., proposed to complete a sequence) during the model’s use would betray its use during training (i.e., it represents a membership inference). As illustrated in Figure 4, the model may be required to memorize contextual information and that used as a mask during the supervised training phase [^6] We discuss context-related memorization in Section 7 and show that its impact on confidentiality is very limited compared to memorizing the words used in masks.. In this example, the fact that the model proposes the word $w4$ in the testing phase means that an adversary could infer that user $u1$’s data was used during training, because this word only appears in this individual’s data in the attacker’s auxiliary knowledge. Note that the indirect identifiers detected from the attacker’s auxiliary information are most likely different from the indirect identifiers from the dataset used to fine-tune the model. This difference can lead to incorrect predictions for the adversary.